### v2

# Learning the Intrinsic Dimensionality of Fermi-Pasta-Ulam-Tsingou Trajectories: A Nonlinear Approach using a Deep Autoencoder Model

Abstract
We address the intrinsic dimensionality (ID) of high-dimensional trajectories, comprising $n_{s}=4\,000\,000$ data points, of the Fermi-Pasta-Ulam-Tsingou (FPUT) $\beta$ model with $N=32$ oscillators. To this end, a deep autoencoder (DAE) is used to infer the ID in the weakly nonlinear regime where energy recurrences are observed ($\beta\lesssim 1$). We find that the trajectories lie on a nonlinear Riemannian manifold of dimension $m^{\ast}=2$ embedded in a $64$-dimensional phase space. By contrast, principal component analysis (PCA) together with the Participation Ratio (PR) method provides only a reasonable upper bound on the ID for each value of $\beta$. Our DAE further reveals that the ID increases to $m^{\ast}=3$ at $\beta=1.1$, coinciding with a symmetry-breaking (SB) phenomenon characteristic of the $\beta$ model, in which additional energy modes with even wave numbers $k=2,4$ become excited. Notably, the SB phenomenon cannot be detected by the linear approach provided by PCA.
I Introduction
Fermi believed that weakly coupled nonlinear systems would inevitably exhibit ergodic behavior, as required to approach thermal equilibrium [22]. As a consequence, in 1953, the first computer simulations of the Fermi-Pasta-Ulam-Tsingou (FPUT) model, a one-dimensional chain of weakly nonlinearly coupled oscillators, were started aiming to study the system’s thermalization, which in turn should have confirmed the validity of the ergodic hypothesis [18, 23, 80, 26, 49, 15]. In contrast, the energy recurrence phenomenon observed from simulations, for which the total energy was shared between a few energy modes only, seemed to contradict the ergodic hypothesis, a fact often referred to as the FPUT paradox.
Since then, the FPUT model has had a huge impact on our understanding of nonlinear dynamics and deterministic chaos. It would be an impossible task to mention all the important contributions made in the last seventy years attempting to explain the peculiar dynamical features of the FPUT chain. Here, we limit ourselves to recalling that the recurrence phenomenon can be linked to the Kolmogorov–Arnol’d–Moser (KAM) theorem [47, 3, 60], the Toda lattice [1, 5], the solitonic behaviour of the Korteweg–de Vries (KdV) equation [82], and q-breathers [19, 20]. Additionally, there exist various theories and methods attempting to explain how the system eventually approaches equilibrium: the existence of a stochastic threshold [39, 53], the FPUT metastability [25, 26, 70], and the spectral-entropy equipartition indicator [52, 26].
Recently, a data-driven approach based on principal component analysis (PCA) was proposed to infer the intrinsic (or effective) dimensionality (ID) $m^{\ast}$ of the high-dimensional trajectories of the FPUT $\beta$ model with $N=32$ oscillators [57]. This work was chiefly motivated by the observation that the ergodic hypothesis requires trajectories to lie on a constant-energy hypersurface of dimension n1n-1 in $\mathbb{R}^{n}$ (n=2Nn=2N[38]. As a result, thermal equilibrium cannot be reached when the trajectories lie on very low-dimensional manifolds, a scenario which aligns with the application of the KAM theorem to the FPUT model in a weakly nonlinear regime [4]. Indeed, according to the KAM theorem, the orbits should be constrained to low-dimensional invariant tori [21, 71].
In this regard, PCA when applied to entire trajectory data consisting of $n_{s}=4\,000\,000$ data points, with the initial condition corresponding to excitation of only the first energy mode ($k=1$), estimates the ID to be either $2$ or $3$ when $\beta\lesssim 1.1$ [57]. These findings strongly suggest non-ergodic behavior, and are consistent with other qualitative observations in the same regime that support the underlying quasi-periodic motion of the FPUT system: the hexagonal patterns shown by Poincaré maps [30], and the energy recurrences observed from the simulations [57].
However, being linear, PCA provides a rather limited approach to the problem at hand [42, 34, 50, 32, 59, 74]. Indeed, it assumes that the data points lie on optimal hyperplanes of a given rank $m$ (or, equivalently, on $m$-dimensional linear subspaces if the data are centered). This assumption is not necessarily satisfied, as shown by a qualitative analysis of early-stage orbits performed using $t$-distributed stochastic neighbor embedding ($t$-SNE), a nonlinear manifold-learning algorithm [78, 45, 46, 13, 57]. More importantly, the inherent limitation of a linear approach was further confirmed by the analysis of trajectory data for $\beta=0.1$ using multi-chart flows, a manifold-learning method [81]. Specifically, this state-of-the-art method suggests that the trajectory lies near a nonlinear Riemannian manifold of dimension $m^{\ast}=2$.
With this in mind, it makes sense to replace PCA with its nonlinear generalization: the autoencoder (AE) [31, 67]. The AE is a type of artificial neural network (ANN) capable of learning the nonlinear structure of data. Specifically, we employ a relatively simple undercomplete deep autoencoder (DAE) to learn the actual nonlinear structure of the trajectory data under scrutiny (see Sec. III). We find that this model can learn and generalise appropriately, yielding negligible reconstruction errors in comparison to those obtained using PCA. From the DAE reconstruction error curves, the corresponding elbow points estimate $m^{\ast}=2$ for all available trajectory data associated with the recurrence phenomenon ($\beta\lesssim 1$) (see Fig. 5).
Finally, the DAE predicts that the intrinsic dimensionality increases by one, i.e., $m^{\ast}=3$, when $\beta=1.1$. We find that this change is associated with the onset of a symmetry-breaking (SB) phenomenon characterizing the $\beta$ model, in which additional energy modes with even wave numbers $k=2,4$ begin to share the total energy of the system. By contrast, this significant variation of ID is completely missed by PCA, for which the estimated dimensionality remains constant as $\beta\to 1.1$.
II FPUT $\beta$ Model
The FPUT model describes a one-dimensional chain of $N$ nonlinear oscillators whose spring constants and masses are set to unity (see Fig. 1). In particular, the $\beta$ ($\beta>0$) model has the following Hamiltonian $H\left(q,p\right)$ where $q=\left(q_{1},\cdots,q_{N}\right)$ and $p=\left(p_{1},\cdots,p_{N}\right)$ [6]
$H\left(q,p\right)={}$ $\frac{1}{2}\sum_{i=1}^{N}p_{i}^{2}+\frac{1}{2}\sum_{i=0}^{N}\left(q_{i+1}-q_{i}\right)^{2}$ (1)
$+\frac{\beta}{4}\sum_{i=0}^{N}\left(q_{i+1}-q_{i}\right)^{4}\,.$
Figure 1: The FPUT chain consists in a discrete system of equal mass points connected by springs. The coordinate $q_{i}$ represents the displacement of $i$-th point from equilibrium.
Accordingly, the equations of motion of the $\beta$ model are the following  [^1] The equations of motion are commonly referred to as the canonical Hamilton equations.
$\begin{dcases}\dot{q}_{i}=p_{i},\\[4.30554pt] \dot{p}_{i}=q_{i+1}+q_{i-1}-2q_{i}+\beta\Big[(q_{i+1}-q_{i})^{3}-(q_{i}-q_{i-1})^{3}\Big].\end{dcases}$ (2)
Note that in Eq. 2 the index $i$ runs from 0 to $N+1$. This is why we assume that the chain has fixed ends, i.e., $q_{0}=q_{N+1}=0$, and hence $p_{0}=p_{N+1}=0$ (Dirichlet boundary conditions).
From a mathematical point of view, the Cauchy problem for the first-order system in Eq. 2 is well posed; that is, the solution exists and is unique because the vector field is analytic, being polynomial, since the number of oscillators $N$ is finite [58]. However, the solution can be obtained only through numerical integration. To this end, we used the velocity Störmer-Verlet scheme, which is a second-order symplectic integrator well suited for Hamiltonian systems such as FPUT chains [33, 6]. Consequently, in the present work, the trajectories were generated with high accuracy using a fixed time step of 0.05 [30].
III A Deep Autoencoder Model
The autoencoder is an ANN trained to reconstruct an input vector $\mathbf{x}\in\mathbb{R}^{n}$ at the output layer, i.e., to produce a reconstruction $\mathbf{\tilde{x}}$, by passing $\mathbf{x}$ through a narrower hidden layer referred to as the latent (or bottleneck) layer. In this layer, the latent (or code) vector $\mathbf{z}\in\mathbb{R}^{m}$ ($1\leq m\ll n$) provides a compressed, low-dimensional, and salient representation of $\mathbf{x}$. This representation is achieved by minimizing the reconstruction error $J_{m}$ (see Eq.  8), or equivalently, the mean squared error (MSE) loss. In general, AEs can be regarded as nonlinear generalizations of PCA: indeed, it can be shown that an autoencoder restricted to be linear is equivalent to PCA [51, 69].
In the present work, we shall employ a deep autoencoder model with five hidden layers $N_{h}=5$, whose architecture with layer widths $[64,32,16,m,16,32,64]$ is schematically depicted in the cartoon shown in Fig. 2. Note that its architecture consists of two sequential networks: the encoder and the decoder. The first maps $\mathbf{x}$ to $\mathbf{z}$, while the second maps $\mathbf{z}$ to $\mathbf{\tilde{x}}$.
Figure 2: Schematic view of a DAE model with $N_{h}=5$ hidden layers. The input and output layers contain $n$ neurons, while the bottleneck layer has $m$ neurons ($m\ll n$). All hidden neurons use ReLU activation function, as indicated by the symbol. The output layer’s neurons employ a linear activation function, which is appropriate for regression.
In the following, we describe all the model hyperparameters, which were chosen based on their excellent reconstruction performance of the input data for latent dimensions $m\geq 2$, together with a preference for simpler models (Ockham’s razor) [40, 29].
The number of neurons (or nodes) per layer, together with the respective activation functions are shown in Table 1. Note that all hidden neurons use ReLU (Rectified Linear Unit) activation function, i.e., $h(y)=\max(0,y)$, for learning nonlinear mappings [62]. By contrast, the output layer employs a linear activation function, which is commonly used for regression tasks.
Layer Type Number of Neurons Activation Function
input layer 64 -
encoder layer 1 32 ReLU
encoder layer 2 16 ReLU
bottleneck layer $m$ ReLU
decoder layer 1 16 ReLU
decoder layer 2 32 ReLU
output layer 64 Linear
Table 1: Number of neurons and corresponding activation functions of the DAE architecture under consideration. The size of the bottleneck layer, $m$, is assumed to satisfy $1\leq m\ll n$, with $n=64$.
The model is trained using the Adam optimizer [44] with a learning rate of $0.001$ and a mini-batch size of $256$, noting that smaller batch sizes (e.g., $32$) do not produce significant differences in performance. The number of training iterations is set to $300$ epochs. However, early stopping (ES), which acts as an implicit regularizer with a patience of $30$ epochs, is chiefly used to avoid unnecessary training, thereby saving significant time and computational resources. Additionally, the ES will retain the model corresponding to the minimum validation error during the optimization of the trainable parameters (weights and biases)  [^2] Note that when the model’s bottleneck has a single neuron, i.e., $m=1$, there are $4,801$ trainable parameters.. We implemented this model using Keras and TensorFlow [28, 12, 2].
In general, especially for neural networks, the training reconstruction error can be very low because the model may fit the training data extremely closely, thereby leading to serious overfitting. As a result, the model may not perform well on unseen data. Accordingly, to ensure that the model generalizes successfully, we split each trajectory dataset into a training set and a test set [35]. The training set contains $70\%$ of the input data, consisting of data points chosen at random, while the test set contains the remaining $30\%$ of the data. Note that $20\%$ of the training set is also used as validation set for monitoring the model’s generalization during training. With this choice, we find that the difference between the training and test reconstruction errors is generally much less $1\%$.
To estimate ID from the reconstruction error curves [10], we need to restrict the latent dimensionality to as low as a single variable ($m=1$). This limitation is inherent to the constraint $m=1$ itself, and holds independently of the specific autoencoder architecture employed as discussed in next Sec. IV. In such a case, it would be unrealistic to expect an autoencoder to capture the complex structure of the $64$-dimensional input data, unless the data lie on an approximately one-dimensional manifold [51], which is not necessarily the case since we estimated ID to be greater than unity ($m^{\ast}=2,3$). Accordingly, this setup will cause a severe information-bottleneck regime, in which the DAE model is expected to perform poorly, but still significantly better than PCA. However, the aforementioned limitation disappears altogether when the latent dimension $m\geq 2$.
It is worth noting that there is considerable room for improvement over the present model, for example by employing learning-rate decay, alternative activation functions such as GeLU, Mish, or SeLU, or a deeper network architecture [24]. To illustrate this, a comparison with two different deeper models with $N_{h}=7$ is briefly addressed in the next section (Sec. IV). Nevertheless, our preference for a simpler model is justified by the reasonably accurate results achieved at a relatively low computational cost.
III.1 Datasets and Data Preprocessing
Each dataset of size $n_{s}=4\,000\,000$ corresponds to a set $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{i},\cdots,\mathbf{x}_{n_{s}}\}$, whose elements $\mathbf{x}_{i}\in\mathbb{R}^{n}$ are data points that belong to a single trajectory obtained from numerical simulations of the FPUT system at a given value of $\beta\in[0.1,1.1]$, sampled at intervals of $\Delta\beta=0.1$ [57].
The set $\mathcal{X}$ is then conveniently arranged in a $n_{s}\times n$ data matrix $X$. However, before feeding it into the DAE model, a data preprocessing step is required, consisting of centering and scaling each variable by its standard deviation, as is customary in PCA [42]. First, centering the data, corresponding to setting the mean of each variable to zero, allows the PCA eigenvalues to be efficiently computed via singular value decomposition (SVD) (see Sec. B). Second, rescaling ensures that each variable, now having unit variance, contributes equally to the analysis regardless of its original scale [42]. This preprocessing step allows for a direct comparison between the DAE results and those obtained from the principal component analysis in Ref. [57].
IV Results and Discussion
In the following, we first compare the performance of different DAEs to demonstrate that the choice of the model described in Section III is reasonable. Second, we focus specifically on the performance of our model. To this end, we study the behaviour of the learning curves obtained during model training on a specific trajectory dataset corresponding to $\beta=0.3$. However, this analysis exemplifies the characteristic behaviour of the model across all the datasets under consideration.
To begin, we compare our relatively simple model with two others featuring a deeper network structure ($N_{h}=7$), with layer widths given by $[64,48,32,16,m,16,32,48,64]$ for bottleneck sizes $m=1$ and $m=2$. The first model differs only in the value of $N_{h}$, while the second introduces additional features: the GeLU (Gaussian Error Linear Unit) activation function and the SOAP (Simultaneous Orthogonalization and Preconditioning) optimizer with an exponential learning rate, replacing ReLU and Adam, respectively. The SOAP optimizer is a recently proposed method that accounts for correlations between parameters, whereas Adam treats them as independent [79]. The DAE incorporating SOAP, GeLU, and an exponential learning rate was suggested and implemented by Froheberg using the JAX library [8] [^3] The learning rate follows an exponential decay schedule with an initial value of $10^{-3}$, decay rate $0.9$, and transition steps of $10^{5}$..
Figure 3: Training and validation MSE loss curves as functions of epoch for bottleneck sizes $m=1,2$ and trajectory data with $\beta=0.3$ from DAE models with (ReLU, $N_{h}=5$; widths $[64,32,16,m,16,32,64]$) and (ReLU, $N_{h}=7$; widths $[64,48,32,16,m,16,32,48,64]$) using the Adam optimizer, and a DAE model with (GeLU, $N_{h}=7$; widths $[64,48,32,16,m,16,32,48,64]$) using the SOAP optimizer. Throughout training, all deep autoencoders employ early stopping with a patience of $30$ epochs.
Figure 4: Training and validation MSE loss curves as functions of epoch for bottleneck sizes $m=1,2$ and trajectory data with $\beta=0.3$ from DAE model with (ReLU, $N_{h}=5$, $[64,32,16,m,16,32,64]$) using Adam optimizer. (Inset) Validation MSE loss curves as functions of epoch for $m=3,4,5$ and same data. Throughout training, the deep autoencoder employs an early stopping with a patience of $30$.
In Fig. 4 the training and validation MSE losses from the different DAEs are shown as functions of epoch for bottleneck sizes $m=1$ and $m=2$. Throughout, ES is set to $30$. The learning curves display significant qualitative and quantitative differences according to the latent dimension under scrutiny. Overall, we do not observe overfitting, as validation curves follow the training curves closely, yielding nearly identical errors. Accordingly, the models generalize well to unseen data.
When $m=1$, regardless of the chosen model, the irregular oscillations observed after a fast initial decay suggest that these autoencoders are approaching their capacity limit, signalling their inability to faithfully reconstruct the high-dimensional data. Overall, they perform similarly, yielding a best MSE of approximately 0.0360.0360.0380.038.
In contrast, when $m=2$, all curves display the expected learning trend: after a fast initial decay they plateau. In this case, the model with GeLU activation outperforms the other two, achieving a best MSE $\approx 4\times 10^{-4}$, roughly one order of magnitude lower than the ReLU-based models ($\approx 45\times 10^{-3}$). However, the learning curves of our model ($N_{h}=5$) are smoother than those of the deeper model ($N_{h}=7$), suggesting that increasing the network depth alone does not guarantee better performance without further architectural or training adjustments.
While all models’ errors are extremely low, especially when compared to those from PCA (see discussion below), the fact that the errors for $m=2$ can be as much as $80$ times smaller than those corresponding to $m=1$ strongly suggests that the intrinsic dimensionality exceeds one.
Next, in Fig. 4 the training and validation MSE losses from our chosen DAE with $N_{h}=5$ and ReLU, are shown as functions of epoch for bottleneck sizes $m=1,2,\cdots,5$. In the main panel, it is now possible to better visualize the curves for $m=1$ and $2$. In both cases, the respective curves exhibit a rapid initial decay followed by plateau, indicating that the model quickly learns the dominant structure of the data before reaching an error floor. Moreover, for each value of $m$, the training and validation curves remain close throughout training, suggesting that the DAE generalizes.
However, when the bottleneck dimension is reduced to $m=1$, the loss saturates at a significantly higher value than for $m=2$. Specifically, considering that early stopping is triggered at around epoch $116$, one finds that the minimum validation MSE loss $\approx 0.036$ is attained at epoch $86$. This is roughly $7$ times larger than the corresponding best MSE loss observed when $m=2$  [^4] The best validation MSE loss is $\approx 0.0049$, attained at epoch $101$ when $m=2$.. Additionally, the model’s capacity limitation due to the severe scalar information bottleneck ($m=1$) is further evidenced by the noisier evolution of the loss curve, including the presence of spikes and late-epoch fluctuations. These observations are indicative of optimization difficulties induced by this regime and cannot be substantially mitigated without increasing the latent dimensionality. By contrast, for $m\geq 2$ the model performs well, yielding substantially smaller reconstruction errors, as further supported by the well-converged validation loss curves for $m=3,4,5$ shown in the inset of Fig. 4.
Next, the DAE reconstruction errors $J_{m}$ as a function of the latent dimension $m$ for a given value of $\beta$, form a curve whose shape can help identifying the trajectory data’s intrinsic dimensionality. Accordingly, the ID can then be inferred from the elbow point of such a curve [10]. This heuristic is based on the assumption that, beyond the elbow, $J_{m}$ no longer decreases significantly as $m$ increases [77]. Note that this approach is inherently subject to a degree of ambiguity and subjectivity, but is still routinely used in the machine learning field, e.g., for manifold reduction and clustering [28]. In the following, we shall automate the identification of the elbow using the Kneedle algorithm (KA) [73], although in the present case the elbow points can be easily identified by visual inspection (which is not the case for strong nonlinearities; see discussion below).
Figure 5: (Top panel) The DAE test reconstruction error curves as functions of the dimension $m$ for data corresponding to $\beta=0.1,0.4,0.7,0.8,0.9,1.0$, and $1.1$. The PCA reconstruction error curve ($\beta=1.1$) is rescaled by a factor of $100$ for proper visualization. (Bottom panel) The DAE reconstruction error curve for data corresponding to $\beta=0.1$.
In Fig. 5 (top panel), some selected representative reconstruction error curves as functions of $m$ are shown. These curves correspond to $\beta=0.1,0.4,0.7,0.8,0.9,1$, and $\beta=1.1$. Additionally, the curve from PCA reconstruction error for $\beta=1.1$ is displayed. The latter has been rescaled by a factor of $100$ to enable visualization alongside the DAE curves, confirming the superior reconstruction performance of the autoencoder. All DAE reconstruction error curves decrease monotonically, except for the one corresponding to $\beta=0.1$, whose actual trend is not visible due to its small magnitude. For this reason, this curve is plotted separately in the bottom panel of Fig. 5 to improve visualization. All curves for which $\beta\leq 1$ exhibit an evident elbow point at $m=2$. Remarkably, this finding agrees with the approach based on multi-chart flows recently proposed by Yu et al. [81]. By contrast, when $\beta=1.1$, the elbow point is shifted by unity, as confirmed by the Kneedle algorithm. Therefore, we can conclude that $m^{\ast}=2$ for $\beta\leq 1$, while $m^{\ast}=3$ for $\beta=1.1$.
Figure 6: Estimated intrinsic dimension $m^{\ast}$ as function of $\beta\in[0.1,1.1]$, using DAE model with Kneedle algorithm (KA) and PCA with the following heuristics: Participation Ratio (PR), Kneedle algorithm (KA), and Kaiser criterion (KC).
Figure 7: Time evolution of the energy $E_{k}$ of the Fourier modes with $k=1,2,3,4,5$ for the $\beta$ model with $N=32$ oscillators. The initial condition corresponds to assigning the total energy $\mathcal{E}_{1}\approx 0.45$ to the first mode $E_{1}$ only.
We note in passing that one would expect a similar approach, based on elbow detection, to be applicable straightforwardly to reconstruction error curves at much larger values of $\beta$, not considered in the present work. Nevertheless, this is not the case, since as $\beta$ increases no evident elbows emerge in the reconstruction error curve. This behaviour is analogous to that observed for PCA reconstruction error curves [57]. To illustrate the above observations, the DAE and PCA curves corresponding to $\beta=1.8$ are shown in Fig. 11 of Sec. C.
That being said, alternative methods will be necessary to ensure a reliable estimation of the ID of trajectory data generated in the strongly nonlinear regime of the FPUT $\beta$ model [83, 81, 7].
We can now compare the intrinsic dimensionalities inferred from the DAE with those previously obtained using PCA [57]. Since PCA admits several alternative heuristics to the Kneedle algorithm, we additionally consider the Kaiser criterion (KC) [41] and the Participation Ratio (PR) [48], both of which are computed directly from the PCA eigenvalues (see Secs. B and B.1 for details on PCA and for the definitions of the heuristics, respectively).
In Fig. 7 we compare the estimated IDs using the DAE model and KA (circles) with those from PCA together with PR (triangles), KA (squares), and KC (inverted triangles), respectively. First, we note that PCA with the KA and KC substantially overestimates the intrinsic dimensionality for $\beta\geq 0.3$, since in such cases $4\leq m^{\ast}\leq 6$. By contrast, PCA together with PR gives $m^{\ast}=2$ and $m^{\ast}=3$ when $\beta\leq 0.3$ and $\beta>0.3$, respectively. These results are in excellent agreement with those obtained from the DAE reconstruction error curves. Nevertheless, we argue that this agreement is purely coincidental, since it is well known that PCA tends to overestimate the intrinsic dimensionality when data have a nonlinear structure, as in the present case. A classical example is the synthetic Swiss roll dataset, for which PCA predicts $m^{\ast}=3$, whereas the data actually lie on a two-dimensional nonlinear manifold ($m^{\ast}=2$[72, 77]. Based on general considerations, PCA should overestimate IDs by at least unity when the data are nonlinear [83]. Consequently, we may conclude that in the present case the linear approach works properly only when $0.3\leq\beta\leq 1$.
However, we note that there is another important difference between the DAE approach and the linear one. Indeed, according to the latter, the estimated ID remains unchanged as $\beta\leq 1.1$, irrespective of the heuristic used, whereas the DAE predicts an increase by unity, i.e., $m^{\ast}=3$. To understand this dimensionality variation, one needs to consider an important feature characterizing the $\beta$ model: the symmetry about its center for initially symmetric excitations [70, 64]. This means that if we initially excite a Fourier mode with an odd wave number, then the energy cannot spread to even-numbered modes. In the present case, the initial condition corresponds to assigning the total energy $\mathcal{E}_{1}\approx 0.45$ to the first mode $E_{1}$ only. As a result, the energy is expected to be shared only by the odd-numbered modes only, in the regime of recurrence. In our case, this occurs when $\beta\leq 1$, as illustrated for instance by the simulation for $\beta=1$ in Fig. 9. In contrast, when $\beta=1.1$, the even-numbered modes $E_{2}$ and $E_{4}$ are also involved in the system’s dynamics, as illustrated in Fig. 7. In such a case, the mode energies $E_{2}$ and $E_{4}$, shown as dotted curves in purple and cyan, respectively, acquire a small amount of energy towards the end of the simulation. However, this energy gain is not attributable to the symplectic integrator numerical error. Indeed, in Sec. A we show that the energy drop of odd modes $E_{1},E_{3}$ and $E_{5}$, which is approximately $6.55\%$ of total energy $\mathcal{E}_{1}$, quantitatively accounts for the energy transferred to even modes $E_{2}$ and $E_{4}$, corresponding approximately $6.02\%$ of $\mathcal{E}_{1}$ (see, in particular Fig. 10). This energy transfer from odd to even modes occurs slowly in a regime where the system has not yet reached thermal equilibrium and cannot be attributed to the characteristic oscillatory time dependence of the symplectic integrator error [16].
That being said, we can relate the change in ID from $m^{\ast}=2$ to $m^{\ast}=3$ to the occurrence of SB in the $\beta$ model, whereby energy begins to flow from odd to even modes. Notably, this important dynamical feature could not be detected using PCA due to the nonlinear structure of the trajectory data. Accordingly, we suspect that the change in ID due to symmetry breaking may imply a substantial difference in the geometric or topological features of the $2$- and $3$-dimensional Riemannian manifolds on which the trajectories lie. This could be properly addressed using suitable tools from topological and geometric data analysis [9, 11, 63, 61, 65, 37, 36].
Figure 8: Two-dimensional DAE embeddings of the entire trajectories comprising $n_{s}=3,000,000$ data points corresponding to $\beta=0.1$. In panels (a) and (b), the embeddings were obtained using ReLU and Linear activation functions in the bottleneck layer, respectively. The reconstruction errors in panels (a) and (b) are the following: $J_{2}\approx 6.15\times 10^{-4}$ and $J_{2}\approx 5.91\times 10^{-4}$ (b), respectively.
Finally, one advantage of using autoencoders is the possibility of visualizing the data and thereby gaining meaningful insights into the dynamics. Accordingly, we compute a two-dimensional latent representation (or embedding) of the entire trajectory, consisting of $n_{s}=3\,000\,000$ data points, generated at very small nonlinearity $\beta=0.1$  [^5] Here we used a smaller number $n_{s}$ of data points to reduce the computational burden.. In particular, we compare the two embeddings obtained when ReLU and Linear activation functions are employed in the bottleneck layer. Panels (a) and (b) of Fig. 8 show the resulting two-dimensional embeddings for the ReLU and Linear cases, respectively. In both cases, the embeddings define compact, ring-like regions with sharp, clearly delineated boundaries.
Note that the embedding in panel (b) confirms that the trajectory data are centered during the preprocessing step. This is not evident in the ReLU case, since the activation function clips negative values to zero. In both cases the two activation functions yield small, comparable reconstruction errors $J_{2}\approx 10^{-4}$.
These two embeddings suggest that the system’s motion is confined to a relatively small region of phase space. As a result, the orbits should lie on a tiny two-dimensional nonlinear Riemannian manifold. In contrast, for larger values of $\beta$, for which the system reaches thermal equilibrium, the corresponding embeddings no longer show a compact ring-like structure, but irregular patterns, consistent with an ergodic exploration of a larger region of phase space. For instance, the characteristic irregular pattern emerging in the strongly nonlinear regime ($\beta=3$) is illustrated in Fig. 12, suggesting that the trajectory is free to explore a much larger region of phase space.
V Conclusion
In this work, we addressed the intrinsic dimensionality of high-dimensional trajectories in the FPUT chain using a deep learning approach. To this end, we employed a deep autoencoder model to analyze the trajectory data. Accordingly, we find that in the weakly nonlinear regime ($\beta\lesssim 1$) the orbits lie on a small two-dimensional nonlinear Riemannian manifold. In contrast, a comparison with a linear approach based on principal component analysis shows that it cannot capture the nonlinear structure of the data, thereby yielding questionable results. Additionally, we showed that PCA is unable to detect crucial dynamical changes, such as the symmetry breaking occurring at $\beta=1.1$. In this case, the autoencoder predicts an increase in the intrinsic dimension by unity, which may be related to significantly different geometric and topological features of the resulting three-dimensional nonlinear manifold.
In future work, it would be highly desirable to validate our findings using alternative approaches that do not rely on heuristics, such as the detection of elbow points [83, 81, 7].
Finally, we conclude noting that the use of manifold learning methods, which rests on the manifold hypothesis, namely, that high-dimensional data lie near or on low-dimensional manifolds [17, 59], is strongly supported by the applicability of the KAM theorem in the weakly nonlinear regime to $\beta$ model [49].
Acknowledgements.
The author thanks Søren Hauberg for advice on correctly training the deep autoencoder, Fabio Frohberg for helpful comments on the manuscript and for allowing the use of his JAX-based deep autoencoder script, and Alberto Maspero for explaining the Cauchy problem for the FPUT system as a first-order differential system. Most of the code execution for this project was performed using Google Colaboratory. The author is also grateful to Google for supporting this research through the Tensor Research Cloud (TRC) program and for providing access to Cloud Tensor Processing Units (TPUs), which accelerated the machine learning experiments.
Data Availability Statement
The data that support the findings of this study are openly available in Zenodo [55, 56].
Appendix A FPUT Initial Condition
In the following, we recall how the “half sine” initial condition of FPUT chain is defined a time $t=0$ [23]. Such a condition corresponds to initially displace the coordinates $q_{i}$ ($i=1,\cdots,32$) according to the formula [30]
$q_{i}\left(0\right)=A\sqrt{\frac{2}{N+1}}\sin\left(\frac{ik\pi}{N+1}\right)\,$, (3)
where the amplitude $A$ and wave number $k$ are set to $A=10$, and $k=1$, respectively. Note that here $q_{0}=q_{N+1}=0$, that is, the chain’s ends are fixed. Additionally, it is assumed that the variables $p_{i}\left(0\right)=0$ with $i=1,\cdots,32$.
Next, the $k$-th energy (Fourier) mode is defined as [30]
$E_{k}=\frac{1}{2}\left[\dot{a}_{k}^{2}+\omega_{k}^{2}a_{k}^{2}\right]\,$, (4)
where $\omega_{k}=2\sin\left(k\pi/2\left(N+1\right)\right)$ is the frequency, $a_{k}$ denotes the normal mode coordinate ($k=1,2,\cdots,N$) which reads [30]
$a_{k}=\sqrt{\frac{2}{N+1}}\sum_{j=0}^{N}q_{j}\sin\left(\frac{jk\pi}{N+1}\right)\,.$ (5)
Note that this large-wavelength (or equivalently low-frequency) initial condition is equivalent to initially give the total energy $\mathcal{E}_{1}\approx 0.45$ to first energy mode $E_{1}$ ($k=1$). In the weakly nonlinear regime ($\beta\lesssim 1$) the recurrence phenomenon occurs. As a result, the total energy is shared by three modes with odd wave numbers $k=1,3,5$, while it cannot spread to modes with even wave numbers due to the parity conservation characterizing the $\beta$ model [64]. This fact is illustrated in Fig. 9 for $\beta=1$.
Figure 9: Recurrence phenomenon. Time evolution of the energy $E_{k}$ of the Fourier modes with $k=1,3,5$ in the FPUT $\beta$ model with $N=32$ coupled oscillators ($\beta=1$). The initial condition corresponds to assigning the total energy $\mathcal{E}_{1}\approx 0.45$ to the first mode $E_{1}$ only.
Next, we show that the energy transfer to modes $E_{2}$ and $E_{4}$ is a direct consequence of the symmetry-breaking feature of the $\beta$ model, rather than an artifact of symplectic integration errors. To this end, we define the following energy fractions: $\left(E_{1}+E_{3}+E_{5}\right)/\mathcal{E}_{1},\quad\left(E_{2}+E_{4}\right)/\mathcal{E}_{1}$, and $\left(\sum_{i=6}^{32}E_{i}\right)/\mathcal{E}_{1}$.
We then examine their time evolution during the system dynamics for $\beta=1.1$ (see Fig. 10). Toward the end of the simulation, the energy of the odd modes $E_{1}$, $E_{3}$, and $E_{5}$ (blue curve) decreases by approximately $6.55\%$, while the even modes $E_{2}$ and $E_{4}$ (brown curve) gain about $6.02\%$ of $\mathcal{E}_{1}$. The difference between these values is therefore small, about $0.53\%$, confirming that an energy redistribution between these modes is taking place.
We also note that the energy of the remaining modes with $k\geq 6$ (grey curve) oscillates within the range $0\%-20\%$, indicating that no systematic energy accumulation occurs in these modes. We therefore conclude that the observed symmetry breaking is not a fictitious effect arising from numerical integration errors.
Figure 10: Time-evolution of the energy fractions (in percentage) $\left(E_{1}+E_{3}+E_{5}\right)/\mathcal{E}_{1}$ (odd modes, blue curve), $\left(E_{2}+E_{4}\right)/\mathcal{E}_{1}$ (even modes, brown color) and $\left(\sum_{i=6}^{32}E_{i}\right)/\mathcal{E}_{1}$ (modes with $k\geq 6$, grey curve).
Appendix B Principal Component Analysis
Let $X^{\ast}$ denote the $n_{s}\times n$ data matrix, after the data preprocessing, which make the variables scale-free each with zero mean, then the sample covariance matrix (or equivalently correlation matrix) $S$ reads [42]
$S=\frac{1}{n_{s}-1}X^{\ast^{T}}X^{\ast}\,.$ (6)
The eigenvalue $\lambda_{i}$ of $S$ correspond to the explained (or preserved) variance along the $i$-th principal component axis. In present work, the sorted eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots\lambda_{i}\geq\cdots\lambda_{n}$ are efficiently computed using the singular value decomposition [76, 75]. Accordingly to SVD, $X^{\ast}=WLV^{T}$ where $W$ and $V$ are two orthogonal matrices, and $L$ is a diagonal matrix [42, 32]. As a result, it is found that [42]
$\lambda_{i}=\frac{1}{n_{s}-1}s_{i}^{2}\,$, (7)
where the singular values $s_{i}$ ($s_{1}^{2}\geq s_{2}^{2}\geq\cdots\geq s_{n}^{2}\geq 0$) are the diagonal entries of $L$, which we computed using scikit-learn library [66].
From the projection perspective PCA orthogonally projects the original data points $\mathbf{x_{j}}$ onto projected data points $\mathbf{\tilde{x}_{j}}$ belonging to a suitable $m$-dimensional linear subspace (the principal subspace), by minimizing the average reconstruction error $J_{m}$ which reads [34]
$J_{m}=\frac{1}{n_{s}}\sum_{j=1}^{n_{s}}\lVert\mathbf{x_{j}}-\mathbf{\tilde{x}_{j}}\rVert_{2}^{2}\,$, (8)
where $\lVert\cdot\rVert_{2}$ stands for the Euclidean norm.
Accordingly, within PCA, Eq. 8 can be rewritten in terms of covariance matrix’s eigenvalues $\lambda_{i}$ as [14]
$J_{m}=\sum_{l=m+1}^{n}\lambda_{l}\,.$ (9)
It is worth noting here that the eigenvectors relative to the eigenvalues $\lambda_{l}$ ($m+1\leq l\leq n$) in Eq. 9 constitute the basis of the orthogonal complement of the principal subspace.
Finally, it is worth noting that within PCA, the term PC score vector (or simply scores) replaces the code vector which sits at the AE’s bottleneck layer [42].
B.1 Heuristics for Determining ID in PCA
Determing the intrinsic dimensionality corresponds to find the optimal number of PCs to retain. To this end, there exist various methods. To our best knowledge, the state-of-art method is represented by the Gavish-Donoho optimal hard threshold, which rests on the Marchenko-Pastur distribution [27, 54]. However, this cannot applied here because the aspect ratio $n/n_{s}\approx 1.6\times 10^{-5}$ of $\tilde{X}$ is too small [57].
Jolliffe modified the Kaiser criterion (also known as the Kaiser-Gutman rule) in factor analysis, making it suitable for PCA [43, 41]. According to Jolliffe’s ansatz, the number of principal components which capture most of the variance, corresponds to those PCs for which $\lambda_{i}\geq 0.7$.
An alternative way to obtain the ID from PCA’s eigenvalues is computing the Participation Ratio $D_{\rm PR}$, which estimates the number of dimensions along which the data is distributed [48, 68]. Accordingly, this quantity reads [68]
$D_{\rm PR}=\frac{\left(\sum\limits_{i=1}^{n}\lambda_{i}\right)^{2}}{\sum\limits_{i=1}^{n}\lambda_{i}^{2}}\,.$ (10)
Eq. 10 can be conventiently expressed in terms of the traces of $S$ (Eq. 6) and $S^{2}$, respectively. In such a case, one finds $D_{\rm PR}=\left(\operatorname{Tr}(S)\right)^{2}/\operatorname{Tr}(S^{2})$. Since $D_{\rm PR}$ is as an estimate of the intrinsic dimensionality of the data, an inherently integer-valued quantity, we round it to the nearest integer.
Appendix C Some DAE Results for $\beta=1.8$ and $\beta=3$
In Fig. 11, the DAE and PCA reconstruction error curves as functions of the latent dimension $m$ are shown for $\beta=1.8$. In this case, most DAE computations were performed using Google TPUs; accordingly, the only modification to the autoencoder architecture was an increase in the batch size from $256$ to $512$ in order to fully exploit the TPUs. However, this change in the autoencoder’s hyperparameters does not affect the overall final result.
Although the curves suggest that the intrinsic dimensionality increases substantially for large $\beta$, no clear elbow points are visible, making their identification highly questionable. Accordingly, by applying KA to the DAE reconstruction curve, and PR, KC, and KA to the PCA reconstruction curve, we obtain the estimates $m^{\ast}=6$ for DAE, and $m^{\ast}=6$, $m^{\ast}=17$, and $m^{\ast}=21$ for PCA, respectively.
Assuming, as discussed above, that the PR provides the most reliable estimate of the intrinsic dimensionality in the linear approach, we conclude that both methods predict the same intrinsic dimension, $m^{\ast}=6$, for data with $\beta=1.8$. Furthermore, DAE aligns with PCA, confirming that the intrinsic dimensionality increases with stronger nonlinearity, as previously reported in Ref. [57].
Figure 11: DAE reconstruction error curve (circle) and PCA reconstruction error curve (triangle), rescaled by a factor of $50$, as functions of latent dimension $m$ for trajectory data corresponding to $\beta=1.8$.
Next, in Fig. 12, the two-dimensional DAE embedding of the entire trajectory with $n_{s}=3\,000\,000$ data points and $\beta=3$ is shown. In contrast to the embeddings obtained for small nonlinearities, i.e., $\beta=0.1$, there is no longer a regular pattern. Additionally, the embedding suggests that the system is free to explore a much larger region of the phase space, consistently with the ergodic hypothesis.
We conclude reporting the intrinsic dimensionality of these trajectory data according to PCA [57]. Accordingly, it is found that $m^{\ast}$ ranges from $10$ (PR) to $37$ (KA, KC).
Figure 12: Two-dimensional DAE embedding of the entire trajectory comprising $n_{s}=3\,000\,000$ data points for $\beta=3$. The embedding is obtained using ReLU activation function in the bottleneck layer. In this case, the reconstruction error is $J_{2}\approx 0.54$.